OpenStudy (amistre64):

Given m and n are in the set of integers, prove directly: If mn=even, then m=even or n=even

OpenStudy (luffingsails):

hmm.. I would start with the thought that since mn is even -- that means there exists a number, k, such that 2 * k = mn.

OpenStudy (amistre64):

good; or in general; mn = 2\(k_1\)

OpenStudy (luffingsails):

More accurately, k is an integer. Therefore, k = mn/2 --> which means mn/2 is an integer.

OpenStudy (anonymous):

An integer is easily defined as even if it can be divided by 2 evenly. Proof: Given all even integers are multiples of 2. Let mn = even , mn = x. Assume that n = 2a+1. 2x (2x+1) = 4x+1 (not even). That was really messy, but you get the idea. It can be easier to prove through contradiction.

OpenStudy (amistre64):

good, the direct proof they say may be a bit difficult to do; but i think it can be done :)

OpenStudy (phi):

It's easy to show if m or n is even, mn/2 is even now show if both m and n are odd, mn/2 has a remainder

OpenStudy (amistre64):

my thought went; \[(2k_1){2k_2+1}=2k_3\] \[4k_1k_2+2k_1=2k_3\] \[2(2(k_1k_2)+k_1)=2k_3\]

myininaya (myininaya):

mn=2k for some integer k mn-2k=0 n(m-2k/n)=0 either n=0 or m-2k/n=0 let's assume n doesn't equal to 0 m=2k/n if n|2k, then m is even and so is n

OpenStudy (amistre64):

dropped some paranthesis there, ....

myininaya (myininaya):

....

OpenStudy (phi):

dividing by n=0 ?

myininaya (myininaya):

?

myininaya (myininaya):

that is why i said assume n does not equal 0

myininaya (myininaya):

i spoke in no particular order

myininaya (myininaya):

its my thinking

OpenStudy (phi):

I don't see how m must even. k could be odd

myininaya (myininaya):

you guys are my digital scratch paper

OpenStudy (amistre64):

\[(2k_1)(2k_2+1)=2k_3\] \[either;\ m=2k_1, or\ n=2k_1\]

myininaya (myininaya):

yes you are right phi if we have 2(3)/2=3 we have odd m

myininaya (myininaya):

assuming n is even (n=2)

myininaya (myininaya):

and some integer k is 3

myininaya (myininaya):

i should had said or instead of and

OpenStudy (phi):

I think a direct proof would go like this both m and n even implies mn is even m even, n odd implies mn is even m odd, n even ibid m odd, n odd mn is odd

OpenStudy (amistre64):

teacher did the last one, proof by contraposition

OpenStudy (phi):

How about arguing m and n can be broken down into their prime factors, and then argue one of m or n must have 2 as a prime factor. We need an if and only if argument